OpenStudy (anonymous):

;If y varies inversely as x, find the constant of variation if y = 5 when x = -1

OpenStudy (lgbasallote):

y varies inversely as x so equation would be y = k/x y = 5 when x = -1 so plug it in 5 = k/-1 solve for k

OpenStudy (diyadiya):

If y varies inversely with x we have this equation: \[y = \frac{1}{k}x\] where k is the constant of variation :D y =5 ,x = -1 now find k

OpenStudy (unklerhaukus):

\[y(x)\propto x^{-1}\] \[y(x)=kx^{-1}=\frac kx\]